OpenStudy (ashking1):

Find an equation for the nth term of a geometric sequence where the second and fifth terms are -8 and 512, respectively. answer choices an = 2 • (-4)^n + 1 an = 2 • 4^n - 1 an = 2 • (-4)^n - 1 an = 2 • 4^n

OpenStudy (johnweldon1993):

You can just do a simple plug and chug process here Check out the 2nd answer choice, lets work with it...you know the second term is -8 so \[\large a_n = 2\times 4^{n-1}\] becomes \[\large a_2 = 2\times 4^{2-1}\] \[\large a_2 = 2\times 4^1\] \[\large a_2 = 2\times 4 = 8\] This is not the correct answer choice since we know the second term SHOULD be -8 not 8 So that eliminates both B and D as answer choices

OpenStudy (ashking1):

would it be c